One side of a triangle is 5 more than the base and the other side is 3 less than twice the base. if the perimeter is 26,find all three sides.

Answer by checkley71(8403)

SIDE1=B+5
SIDE2=2B-3
BASE=B
26=B+B+5+2B-3
26=4B+2
4B=26-2
4B=24
B=24/4
B=6 ANSWER FOR THE BASE.
6+5=11 FOR SIDE1.
2*6-3=12-3=9 FOR SIDE2.
PROOF:
26=6+11+9
26=26